### v4

@@ 3.1 The Randomized Nonlocal Operator
    For any $\phi_{1},\phi_{2}\in\mathscr{F}$, $\left\|\mathcal{M}^{\lambda}\phi_{1}-\mathcal{M}^{\lambda}\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}\leq\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}$.
@@
The next results concern with $\mathcal{M}^{\lambda}$ being applied to more restricted families of test functions, in particular, Lipschitz test functions. This additional restriction is also in alignment with the Lipschitz value function introduced in Definition 2.4.
@@ 4.2.1 Connection with Randomization via Controlled Compound Poisson Process
Inspired by denkert2025control , we embed the impulse mechanism via a controlled compound Poisson process. Consider $\pi=\{\pi_{t}\}_{t\geq 0}$ such that $\pi_{t}\geq 0$ for all $t\geq 0$, and let $N^{\pi}=\{N^{\pi}_{t}\}_{t\geq 0}$ be a nonhomogeneous Poisson process with intensity $\pi$ such that, conditioning on $\pi$, $N^{\pi}\perp W$. Let $\eta=\{\eta_{n}\}_{n=1}^{\infty}$ be a sequence of independent jumps such that $\eta\perp(W,N^{\pi})$. Define the following jump process $Y^{\pi,\bm{\mu}}=\{Y^{\pi,\bm{\mu}}_{t}\}_{t\geq 1}$ such that
@@ 5 Solving Randomized Impulse Control Problem via an Iterative Approach
In this section, we provide a constructive proof for the existence of solution to the randomized impulse control problem specified in Definition 2.4 under a fixed $\bm{\lambda}\in(0,+\infty)^{2}$. Its building block is an iterative approach mirroring the classical Howard iteration howard1960dynamic . The construction of the iterates consists of two key components, namely the “policy evaluation” and the “policy improvement”. Tailored to our randomized impulse control problem, these two components are achieved through the compound optimal stopping operator $\mathcal{T}^{\bm{\lambda}}$ defined in (randomized). We show that this iterative approach leads to a limiting function, which is indeed a value function under Definition 2.4.
@@ 5.1 Construction of the iterates
subject to (2.1) and (2.12) with $p_{0}=1$. Then we have the following property for the iterates $\psi^{\bm{\lambda},n}$’s.
@@ 6 Convergence of the Randomized Problem to the Classical Problem
$\psi^{n}(x)=\mathcal{T}\psi^{n-1}(x)=\inf_{\tau}\mathbb{E}_{x}\left[\int_{0}^{\tau}e^{-rt}f(X_{t})\mathrm{d}t+\mathcal{M}\psi^{n-1}(X_{t})\right],$ (6.1)
subject to (2.1).
Equivalently, one can show by classical results bensoussan2011applications ; pham2009continuous that when $n\in\mathbb{N}^{+}$, $\psi^{n}$’s are the unique viscosity solution to the corresponding HJB equation:
@@
The first condition is a common convention in the literature on impulse control problems reisinger2019penalty ; reisinger2020error . The second condition is satisfied in many practical settings. For instance, both our example in Section 7.2 and the numerical study in reisinger2020error meet this requirement.
@@
Proposition 6.2 ((reisinger2019penalty, , Theorem 3.4) and (reisinger2020error, , Proposition 4.1))
@@
Given the similarity in the problem settings, the proof of Proposition 6.2 is essentially the same as that of (reisinger2020error, , Proposition 4.1), therefore omitted.
@@
In fact, one can show by going through the same line of proof as former work (pham2009continuous, , Section 3,5) that for any $n\in\mathbb{N}$, the newly-introduced semi-randomized solution $\tilde{\psi}^{n}$ is the unique viscosity solution to (6.4). Moreover, we have the following propositions.
@@
This follows directly from (pham2009continuous, , Section 3,5) and bensoussan2011applications since the obstacle function $\mathcal{M}^{\lambda_{2}}\psi^{n}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$, therefore the proof is omitted.
@@ 7.2 A Temporal-Difference Error Method
Now, we present a reinforcement learning algorithm for the randomized impulse control problem with a TD loop. In particular, we use a TD scheme to estimate the “policy evaluation” step inside the randomized impulse–control framework.
@@
2:for $n=0,1,\dots,N-1$ do
3:  $\psi^{n}(\cdot)\leftarrow\psi(\cdot;\theta_{0}^{n})$.
4:  for $b=0,1,\dots,B-1$ do
5:   Sample $X_{0}^{b}\sim\rho$.
6:   Simulate $\{X_{t_{i}}^{b}\}_{i=0}^{M}$ via the controlled SDE drift/diffusion.
7:   For all $i$, compute $\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t_{i}}^{b})$ (Monte Carlo method).
8:  end for
@@ 7.4.1 Comparison between randomized value functions and their classical counterpart
In Figure 1, we present the comparison results between the randomized and the classical value functions, as $\bm{\lambda}$ varies in $\Lambda$. We plot the corresponding value functions in Figure 1a, where the randomized value functions $\psi^{N}$ are the final outputs of Algorithm 1 and the classical value function $\psi$ is the semi-explicit analytical solution as in constantinides1978existence ; guo2009smooth . It can be observed that as $\bm{\lambda}$ tends to $(0,0)$, the randomized solution approaches the classical benchmark $\psi$, which numerically confirms the theoretical limit $\psi^{\bm{\lambda}}\xrightarrow[]{\|\bm{\lambda}\|_{\infty}\to 0}\psi$ as proved in Theorem 6.9.
@@ 7.4.2 Sensitivity Analysis
For the classical impulse control in constantinides1978existence , the size of the continuation region $|\mathfrak{C}|=O\left(\sigma^{\frac{2}{3}}+\sigma\right)$, and the jump distance $|\xi^{*}(x)|=O\left(\sigma^{\frac{2}{3}}\right)$ for $x$ on the boundary of the continuation region. That is, as $\sigma$ increases, the controller intervenes less frequently, but once hitting the action boundary from inside the continuation region, the jump size is larger. Intuitively, with a larger $\sigma$ value, the controller may make use of the extra randomness when deciding whether to intervene, but may need to compensate for the extra uncertainty by jumping further into the continuation region during each intervention. Consequently, by accumulating more running cost and paying higher cost of intervention, the corresponding value function tends to increase along with $\sigma$. We can observe similar trends under the randomized setting. For instance, in Figure 2a, we observe that the randomized value function monotonically increases as $\sigma$ increases. In Figure 2b, we visualize the optimal intensities $\pi^{*}$ for the survival processes under different values of $\sigma$. As $\sigma$ increases, the flat region where $\pi^{*}$ remains close to 0 (i.e., it is more likely for the controller to wait) keeps expanding. This observation coincides with the enlarging continuation region in the classical problem.
@@ A.1 Proofs in Section 3
Without loss of generality, assume $\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}<\infty$. For any $x\in\mathbb{R}$ and any $\epsilon>0$, $\exists\mu_{\epsilon}\in\mathcal{P}\left(\mathbb{R}\right)$ such that
@@
$\mathcal{M}^{\lambda}\phi(x)=\inf_{\rho\geq 0,\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\right]}\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\left(\phi(x+\zeta)+l(\zeta)\right)+\lambda\rho(\zeta)\log\rho(\zeta)\right],\quad\forall x\in\mathbb{R}.$ (A.1)
Notice that the mapping $(0,+\infty)\ni r\mapsto cr+\lambda r\log{r}$ is convex for any constant $c\in\mathbb{R}$. For any $x,\zeta\in\mathbb{R}$, denote
@@
From (guo2009smooth, , Lemma 2.2-(3)), we know that $\mathcal{M}\phi\in\mathscr{L}$. Then the conclusion follows from Dini’s theorem.
@@
Sending $\epsilon\to 0$ yields closes the proof.
@@
Together with the standing Assumptions 2.1 and 2.3, we can apply existence theorems for quasi-linear elliptic equations (see (gilbarg2013elliptic, , Theorem 15.10.) for instance) as well as the comparison principle (see (gilbarg2013elliptic, , Theorem 10.1) for instance) to conclude that (A.4) admits a unique classical solution $v_{N}\in\mathcal{C}_{loc}^{2,\alpha}\cap\mathcal{C}$ for any $\alpha\in(0,1)$. Notice that on $\bar{\Omega}_{N}\times\mathbb{R}\times\mathbb{R}$,
@@
Notice that the constants $C_{1}=\frac{\sigma_{0}^{2}}{2}\cdot\sup_{x\in\Omega_{N}}b(x)$ and $C_{2}=\frac{\sigma_{0}^{2}}{2}[\lambda e^{-\frac{l_{\phi}}{\lambda}}+\sup_{x\in\Omega_{N}}f(x)]\geq 0$ depend only on $(L,\sigma_{0},L_{f},l_{\phi},\lambda,N)$. By maximum principle (gilbarg2013elliptic, , Theorem 10.3), $v_{N}$ is bounded uniformly for all $M>0$. Therefore, given $M$ sufficiently large, $v_{N}$ satisfies (3.5) on $\Omega_{N}$.
Now, choosing any $q>1$ and any $\Gamma\Subset\Omega_{N}$ and applying interior estimate (for example, (gilbarg2013elliptic, , Theorem 9.11.)), we have
@@
with some constant $c^{\prime}>0$. By Sobolev embedding theorem (see (Evans1998, , Section 5, Theorem 6) for instance), we have that

### v5

@@ 3.1 The Randomized Nonlocal Operator
    For any $\phi_{1},\phi_{2}\in\mathscr{F}$ with $\phi_{1}-\phi_{2}\in L^{\infty}(\mathbb{R})$, $\left\|\mathcal{M}^{\lambda}\phi_{1}-\mathcal{M}^{\lambda}\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}\leq\left\|\phi_{1}-\phi_{2}\right\|_{L^{\infty}(\mathbb{R})}$.
@@
The next results concern with $\mathcal{M}^{\lambda}$ being applied to more restricted families of test functions, in particular, Lipschitz test functions. This additional restriction is also in alignment with the Lipschitz value function introduced in Definition 2.4. Moreover, a closed-form expression for our randomized operator is provided.
@@ 4.2.1 Connection with Randomization via Controlled Compound Poisson Process
Inspired by Denkert et al. (2025), we embed the impulse mechanism via a controlled compound Poisson process. Consider $\pi=\{\pi_{t}\}_{t\geq 0}$ such that $\pi_{t}\geq 0$ for all $t\geq 0$, and let $N^{\pi}=\{N^{\pi}_{t}\}_{t\geq 0}$ be a nonhomogeneous Poisson process with intensity $\pi$ such that, conditioning on $\pi$, $N^{\pi}\perp W$. Let $\eta=\{\eta_{n}\}_{n=1}^{\infty}$ be a sequence of independent jumps such that $\eta\perp(W,N^{\pi})$. Define the following jump process $Y^{\pi,\bm{\mu}}=\{Y^{\pi,\bm{\mu}}_{t}\}_{t\geq 1}$ such that
@@ 5 Solving Randomized Impulse Control Problem via an Iterative Approach
In this section, we provide a constructive proof for the existence of solution to the randomized impulse control problem specified in Definition 2.4 under a fixed $\bm{\lambda}\in(0,+\infty)^{2}$. Its building block is an iterative approach mirroring the classical Howard iteration Howard (1960). The construction of the iterates consists of two key components, namely the “policy evaluation” and the “policy improvement”. Tailored to our randomized impulse control problem, these two components are achieved through the compound optimal stopping operator $\mathcal{T}^{\bm{\lambda}}$ defined in (randomized). We show that this iterative approach leads to a limiting function, which is indeed a value function under Definition 2.4. The proofs are deferred to Appendix A.2.
@@ 5.1 Construction of the iterates
subject to (2.1) and (2.12) with $p_{0}=1$.
Intuitively, the sequence $\{\psi^{\bm{\lambda},n}\}_{n\geq 0}$ is constructed via the compound randomized optimal stopping operator $\mathcal{T}^{\bm{\lambda}}$. At each step, $\psi^{\bm{\lambda},n}$ and $\pi^{n}$ respectively represent the value function and the optimal stopping rule for the randomized optimal stopping problem with terminal cost $\mathcal{M}^{\lambda_{2}}\psi^{\bm{\lambda},n-1}$. Since an optimal stopping problem only solves for the optimal intervention time, not for the jump size, the policy iteration naturally acts on the survival rate $\pi^{n}$. By contrast, the optimal jump-size distribution $\mu$ is already available in closed form (3.1) thanks to the regularization via KL divergence, and hence no additional inner optimization over distributions is required.
Furthermore, we have the following property for the iterates $\psi^{\bm{\lambda},n}$’s.
@@ 6 Convergence of the Randomized Problem to the Classical Problem
$\psi^{n}(x)=\mathcal{T}\psi^{n-1}(x)=\inf_{\tau}\mathbb{E}_{x}\left[\int_{0}^{\tau}e^{-rt}f(X_{t})\mathrm{d}t+e^{-r\tau}\mathcal{M}\psi^{n-1}(X_{\tau})\right]$}, (6.1)
subject to (2.1) and $X_{0}=x$.
Equivalently, one can show by classical results Bensoussan and Lions (2011); Pham (2009) that when $n\in\mathbb{N}^{+}$, $\psi^{n}$’s are the unique viscosity solution to the corresponding HJB equation:
@@
The first condition is a common convention in the literature on impulse control problems Reisinger and Zhang (2019, 2020). The second condition is satisfied in many practical settings. For instance, both our example in Section 7.2 and the numerical study in Reisinger and Zhang (2020) meet this requirement.
@@
Proposition 6.2 ((Reisinger and Zhang, 2019, Theorem 3.4) and (Reisinger and Zhang, 2020, Proposition 4.1))
@@
Given the similarity in the problem settings, the proof of Proposition 6.2 is essentially the same as that of (Reisinger and Zhang, 2020, Proposition 4.1), therefore omitted.
@@
In fact, one can show by going through the same line of proof as former work (Pham, 2009, Section 3,5) that for any $n\in\mathbb{N}$, the newly-introduced semi-randomized solution $\tilde{\psi}^{n}$ is the unique viscosity solution to (6.4). Moreover, we have the following propositions.
@@
This follows directly from (Pham, 2009, Section 3,5) and Bensoussan and Lions (2011) since the obstacle function $\mathcal{M}^{\lambda_{2}}\psi^{n}\in\mathscr{L}_{\frac{L_{f}}{r-G}}$, therefore the proof is omitted.
@@ 7.2 A Temporal-Difference Error Method
Now, we present a reinforcement learning algorithm for the randomized impulse control problem with a TD loop. In particular, we use a TD scheme to estimate the “policy evaluation” step inside the randomized impulse–control framework. Our code is publicly available at https://github.com/Zhouhao-Yang/A-Two-fold-Randomization-Framework-for-Impulse-Control-Problems.
Throughout this section we work in a simulator-based setting: we assume the structural form of the uncontrolled dynamics and costs, but we do not assume that the drift $b$, volatility $\sigma$, or cost functions $f$ and $l$ are known in closed form. Instead, the algorithm only interacts with the environment through simulated trajectories and pointwise cost evaluations. All of our updates are therefore expressed in terms of these simulated paths and cost samples, without requiring explicit knowledge of $b$, $\sigma$, $f$, or $l$.
@@
2:for $b=0,1,\dots,B-1$ do
3:  Sample $X_{0}^{b}\sim\rho$.
4:  Simulate $\{X_{t_{i}}^{b}\}_{i=0}^{M}$ via the controlled SDE drift/diffusion.
5:end for
6:for $n=0,1,\dots,N-1$ do
7:  $\psi^{n}(\cdot)\leftarrow\psi(\cdot;\theta_{0}^{n})$.
8:  For all $i$, compute $\mathcal{M}^{\lambda_{2}}\psi^{n}(X_{t_{i}}^{b})$ (Monte Carlo method).
@@ 7.3 Experimental Setting
Classical model-based baseline
For comparison, we also implement a classical, model-based baseline on the same example. In this baseline we assume that only discrete observations of the uncontrolled diffusion $\{X_{t_{i}}^{b}\}$ are available, and we first estimate the drift and volatility via the standard moment estimators
$\widehat{b}=\frac{1}{N\Delta t}\sum_{i=0}^{N-1}(X_{t_{i+1}}-X_{t_{i}}),\qquad\widehat{\sigma}^{2}=\frac{1}{N\Delta t}\sum_{i=0}^{N-1}(X_{t_{i+1}}-X_{t_{i}})^{2}.$
[…]
$\min\bigl\{\hat{\mathcal{L}}\psi^{n}-r\psi^{n}+f,\;\mathcal{M}\psi^{n-1}-\psi^{n}\bigr\}=0,\qquad n\in\mathbb{N}^{+},$
[…]
@@ 7.4.1 Comparison between randomized value functions and their classical counterpart
In Figure 1, we present the comparison results between the randomized and the classical value functions, as $\bm{\lambda}$ varies in $\Lambda$. We plot the corresponding value functions in Figure 1a, where the randomized value functions $\psi^{N}$ are the final outputs of Algorithm 1 and the classical value function $\psi$ is the semi-explicit analytical solution as in Constantinides and Richard (1978); Guo and Wu (2009). It can be observed that as $\bm{\lambda}$ tends to $(0,0)$, the randomized solution approaches the classical benchmark $\psi$, which numerically confirms the theoretical limit $\psi^{\bm{\lambda}}\xrightarrow[]{\|\bm{\lambda}\|_{\infty}\to 0}\psi$ as proved in Theorem 6.9.
In addition to the analytic classical value function $\psi$, fig. 1a also displays a classical, model-based baseline (the curve labeled as “Baseline”). The baseline curve exhibits a noticeable upward bias relative to the analytic classical solution, both in level and in the location of the continuation region. This bias can be traced to three main sources: (i) the inaccuracy in estimating the drift $b$ from a finite number of sample paths; (ii) the inaccuracy of the HJB–QVI solver as we could only solve for a finite region; and (iii) the lack of robust convergence of the classical fixed-point policy iteration. By contrast, our method circumvent these problems, and the randomized value functions learned by algorithm 1 remain much closer to the analytic solution across the state space.
@@ 7.4.2 Sensitivity Analysis
For the classical impulse control in Constantinides and Richard (1978), the size of the continuation region $|\mathfrak{C}|=O\left(\sigma^{\frac{2}{3}}+\sigma\right)$, and the jump distance $|\xi^{*}(x)|=O\left(\sigma^{\frac{2}{3}}\right)$ for $x$ on the boundary of the continuation region. That is, as $\sigma$ increases, the controller intervenes less frequently, but once hitting the action boundary from inside the continuation region, the jump size is larger. Intuitively, with a larger $\sigma$ value, the controller may make use of the extra randomness when deciding whether to intervene, but may need to compensate for the extra uncertainty by jumping further into the continuation region during each intervention. Consequently, by accumulating more running cost and paying higher cost of intervention, the corresponding value function tends to increase along with $\sigma$. We can observe similar trends under the randomized setting. For instance, in Figure 2a, we observe that the randomized value function monotonically increases as $\sigma$ increases. In Figure 2b, we visualize the optimal intensities $\pi^{*}$ for the survival processes under different values of $\sigma$. As $\sigma$ increases, the flat region where $\pi^{*}$ remains close to 0 (i.e., it is more likely for the controller to wait) keeps expanding. This observation coincides with the enlarging continuation region in the classical problem.
@@ 8 Conclusion
Limitations and Future Work
Our two-fold randomization framework and the associated RL algorithm are expected to extend to higher dimensions, while the main additional challenge concerns with the analysis of the nonlocal operator. In higher dimensions, the uncontrolled state process satisfies $X_{t}\in\mathbb{R}^{d}$ and the intervention distribution becomes a law on $\mathbb{R}^{d}\times[0,\infty)$, so that the jump size and direction, as well as the intervention time, are randomized. The corresponding randomized HJB equation has the same structural form. Under similar structural assumptions on the drift, diffusion, and cost (i.e., Lipschitz continuity and at most linear growth), we expect well-posedness of the randomized problem (i.e., existence, uniqueness, and $C^{2}$-regularity) to carry over to $d>1$.
One foreseeable additional difficulty arises in the convergence result (Theorem 6.9), where we show that the value function of the randomized problem converges to that of the classical impulse-control problem as the randomization parameter vanishes. Our current proof relies on the one-dimensional setting, in particular the structure of the randomized nonlocal operator and the local estimates in Lemma 3.2, which do not immediately extend to $\mathbb{R}^{d}$. In higher dimension, one must control both jump directions and geometry-induced effects in the nonlocal term, and hence extending Lemma 3.2 and the subsequent convergence argument to higher dimensions would require more delicate estimates for multidimensional operators.
@@ A.1 Proofs in Section 3
For any $x\in\mathbb{R}$ and any $\epsilon>0$, $\exists\mu_{\epsilon}\in\mathcal{P}\left(\mathbb{R}\right)$ such that
@@
$\mathcal{M}^{\lambda}\phi(x)=\inf_{\rho\geq 0,{\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\right]=1}}\mathbb{E}_{\zeta\sim\Phi_{x}}\left[\rho(\zeta)\left(\phi(x+\zeta)+l(\zeta)\right)+\lambda\rho(\zeta)\log\rho(\zeta)\right],\ \ \forall x\in\mathbb{R}.$ (A.1)
Let us find the infimum of (A.1). Notice that the mapping $(0,+\infty)\ni r\mapsto cr+\lambda r\log{r}$ is convex for any constant $c\in\mathbb{R}$, whose minimum is attained at $r^{*}=\exp(-\tfrac{c+\lambda}{\lambda})$.
For any $x,\zeta\in\mathbb{R}$, denote
@@
$\lambda_{0}D_{\mathrm{KL}}(\Phi_{\sigma}\|\Phi_{x})=\lambda_{0}\,\left(-\log\sigma+\frac{\sigma^{2}+(c+x)^{2}-1}{2}\right)\leq\frac{u-\ell}{2}.$}
And hence,
@@
From (Guo and Wu, 2009, Lemma 2.2-(3)), we know that $\mathcal{M}\phi\in\mathscr{L}$. Then the conclusion follows from Dini’s theorem.
@@
$=$ $\max\left\{\frac{L_{f}}{r-G},L_{\phi}\right\}|x-y|\mathbb{E}\left[\int_{0}^{\infty}\mathrm{d}\left(e^{-(r-G)t}p_{t}\right)\right]$
@@
Sending $\epsilon\to 0$ closes the proof.
@@
Together with the standing Assumptions 2.1 and 2.3, we can apply existence theorems for quasi-linear elliptic equations (see (Gilbarg and Trudinger, 2013, Theorem 15.10.) for instance) as well as the comparison principle (see (Gilbarg and Trudinger, 2013, Theorem 10.1) for instance) to conclude that (A.4) admits a unique classical solution $v_{N}\in\mathcal{C}_{loc}^{2,\alpha}\cap\mathcal{C}$ for any $\alpha\in(0,1)$. Notice that on $\bar{\Omega}_{N}\times\mathbb{R}\times\mathbb{R}$,
@@
Notice that the constants $C_{1}=\frac{\sigma_{0}^{2}}{2}\cdot\sup_{x\in\Omega_{N}}b(x)$ and $C_{2}=\frac{\sigma_{0}^{2}}{2}[\lambda e^{-\frac{l_{\phi}}{\lambda}}+\sup_{x\in\Omega_{N}}f(x)]\geq 0$ depend only on $(L,\sigma_{0},L_{f},l_{\phi},\lambda,N)$. By maximum principle (Gilbarg and Trudinger, 2013, Theorem 10.3), $v_{N}$ is bounded uniformly for all $M>0$. Therefore, given $M$ sufficiently large, $v_{N}$ satisfies (3.5) on $\Omega_{N}$.
Now, choosing any $q>1$ and any $\Gamma\Subset\Omega_{N}$ and applying interior estimate (for example, (Gilbarg and Trudinger, 2013, Theorem 9.11.)), we have
@@
with some constant $c^{\prime}>0$. By Sobolev embedding theorem (see (Evans, 1998, Section 5, Theorem 6) for instance), we have that